OpenStudy (surfkitty05):

HELP, please! (Show work) A gas company has the following rate schedule for natural gas usage in single-family residences: Monthly service charge $8.80 Per therm service charges: 1st 25 therms - $0.6686/therm Over 25 therms - $0.85870/therm What is the charge for using 25 therms in one month. What is the charge for using 45 therms in one month. Construct a function that gives the monthly charge C for x therms of gas.

OpenStudy (surfkitty05):

@misty1212

OpenStudy (surfkitty05):

would you be ok helping me for awhile XD

OpenStudy (surfkitty05):

@inkyvoyd maybe you could help me

OpenStudy (ineedhelplz):

Monthly service charge: $8.80 Per therm service charge 1st 25 therms: $0.6686/therm Over 25 therms: $0.85870/therm therefore for the first 25 therms in a month, 8.80 + 0.6686(25) = ? and for 45 therms in a month, 8.80 + 0.6686(25) + 0.85870(45-25) = ? For x therms in a month, Between zero and 25: f(x) = 8.80 + 0.6686x for 0 ≤ x ≤ 25 Greater than 25: f(x) = 8.80 + 0.6686(25) + 0.85870(x-25) for x > 25 Just simplify this function.